Find all negative odd integers that satisfy the following inequality
-3x+1=<17

Answer by sdmmadam@yahoo.com(530)

This problem I have already worked out under someother number.
Please verify.
Anyway here is the answer
Find all negative odd integers that satisfy the following inequality
-3x+1=<17 ----(1)
-3x=<17-1 (adding minus 1 to both the sides)
-3x=<16
(-3)x/(-3)>= 16/(-3)(multiplication or division by a negative quantity alters the inequality)
x>= (-16/3)which means x is to the right of -5 and (1/3)
and given that x is a negative odd integer.
Therefore our set is {-5,-3,-1}